# ANLS* - A Universal Document Processing Metric for Generative Large Language Models

### v5

@@ Abstract
This paper introduces a new metric for generative models called ANLS* for evaluating a wide variety of tasks, including information extraction and classification tasks. The ANLS* metric extends existing ANLS metrics as a drop-in-replacement and is still compatible with previously reported ANLS scores. An evaluation of 7 different datasets, 6 different GLLMs and 3 different prompting methods using the ANLS* metric is also provided, demonstrating the importance of the proposed metric.
We also benchmark a novel approach to generate prompts for documents, called SFT, against other prompting techniques such as LATIN. In 27 out of 35 cases, SFT outperforms other techniques and improves the state-of-the-art, sometimes by as much as $18$ percentage points.
@@ 4.2 Experimental Evaluation
Note that special prompting techniques were required for mistral-large as well as claude-3 in order to get decent values. We refer to the source code for more details. The results are shown in Table 5. For gemini-pro we set the safety filter to ”High” to not measure false positives of the safety filter.
Table 3: ANLS* score for different GLLMs and Datasets.
@@
Dataset Method gpt-4-turbo gpt-4-vision gemini-1.5-pro mistral-large claude-3
DocVQA Simple 0.589 0.759 0.647 0.445 0.768
Latin Prompting 0.631 - 0.668 0.447 0.762
SFT (Ours) 0.770 - 0.737 0.648 0.831
MPDocVQA Simple 0.606 0.708 0.627 0.364 0.636
Latin Prompting 0.656 - 0.726 0.335 0.438
SFT (Ours) 0.760 - 0.794 0.476 0.575
Kleister Charity Simple 0.708 0.751 0.764 0.652 0.800
Latin Prompting 0.728 - 0.763 0.576 0.787
SFT (Ours) 0.733 - 0.779 0.657 0.786
Kleister NDA Simple 0.633 0.664 0.722 0.637 0.673
Latin Prompting 0.643 - 0.707 0.624 0.670
SFT (Ours) 0.659 - 0.715 0.641 0.677
SROIE Simple 0.849 0.834 0.594 0.855 0.933
Latin Prompting 0.847 - 0.623 0.863 0.926
SFT (Ours) 0.890 - 0.623 0.905 0.949
VRDU AD Buy Simple 0.483 0.640 0.645 0.386 0.577
Latin Prompting 0.506 - 0.686 0.435 0.608
SFT (Ours) 0.712 - 0.765 0.594 0.633
VRDU Registration Simple 0.609 0.665 0.730 0.579 0.685
Latin Prompting 0.614 - 0.749 0.587 0.715
SFT (Ours) 0.590 - 0.780 0.639 0.705

### v6

@@ Abstract
This paper introduces a new metric for generative models called ANLS* for evaluating a wide variety of tasks, including information extraction and classification tasks. The ANLS* metric extends existing ANLS metrics as a drop-in-replacement and is still compatible with previously reported ANLS scores. An evaluation of 7 different datasets, and more than 10 different GLLMs together with 3 different prompting methods using the ANLS* metric is also provided, demonstrating the importance of the proposed metric.
We also benchmark a novel approach to generate prompts for documents, called SFT, against other prompting techniques such as LATIN. In 6 out of 7 cases, SFT outperforms other techniques and improves the state-of-the-art, sometimes by as much as $10$ percentage points.
@@ 4.2 Experimental Evaluation
Note that special prompting techniques were required for mistral-large as well as claude-3 in order to get decent values. We refer to the source code for more details. The results are shown in Table 3. More detailed results with intermediate versions of each model are given in the appendix. For gemini-pro we set the safety filter to ”High” to not measure false positives of the safety filter.
Table 3: ANLS* score for different GLLMs and Datasets. For each model the latest version is reported. Scores for older versions can be found in the appendix.
@@
Dataset Method gpt-4-o gpt-4-vision gemini-1.5-pro mistral-large claude-35
DocVQA Simple 0.656 0.759 0.647 0.445 0.653
Latin Prompting 0.718 - 0.668 0.447 0.739
SFT (Ours) 0.747 - 0.737 0.648 0.769
MPDocVQA Simple 0.686 0.708 0.627 0.364 0.581
Latin Prompting 0.718 - 0.726 0.335 0.561
SFT (Ours) 0.762 - 0.794 0.476 0.582
Kleister Charity Simple 0.711 0.751 0.764 0.652 0.774
Latin Prompting 0.721 - 0.763 0.576 0.786
SFT (Ours) 0.737 - 0.779 0.657 0.773
Kleister NDA Simple 0.649 0.664 0.722 0.637 0.623
Latin Prompting 0.646 - 0.707 0.624 0.626
SFT (Ours) 0.655 - 0.715 0.641 0.635
SROIE Simple 0.887 0.834 0.594 0.855 0.922
Latin Prompting 0.863 - 0.623 0.863 0.89
SFT (Ours) 0.897 - 0.623 0.905 0.917
VRDU AD Buy Simple 0.596 0.640 0.645 0.386 0.569
Latin Prompting 0.655 - 0.686 0.435 0.609
SFT (Ours) 0.753 - 0.765 0.594 0.612
VRDU Registration Simple 0.605 0.665 0.730 0.579 0.663
Latin Prompting 0.611 - 0.749 0.587 0.664
SFT (Ours) 0.605 - 0.780 0.639 0.67
@@ Appendix B Results Gemini
Appendix C Results Claude
In this appendix, a detailed evaluation of different versions of Claude is shown.
Table 6: ANLS* score for different versions of Claude.
Dataset Method claude-3-opus-20240229 claude-3-5-sonnet-20240620
DocVQA Simple 0.768 0.653
Latin Prompting 0.762 0.739
SFT (Ours) 0.831 0.769
MPDocVQA Simple 0.636 0.581
Latin Prompting 0.438 0.561
SFT (Ours) 0.575 0.582
Kleister Charity Simple 0.800 0.774
Latin Prompting 0.787 0.786
SFT (Ours) 0.786 0.773
Kleister NDA Simple 0.673 0.623
Latin Prompting 0.670 0.626
SFT (Ours) 0.677 0.635
SROIE Simple 0.933 0.922
Latin Prompting 0.926 0.89
SFT (Ours) 0.949 0.917
VRDU AD Buy Simple 0.577 0.569
Latin Prompting 0.608 0.609
SFT (Ours) 0.633 0.612
VRDU Registration Simple 0.685 0.663
Latin Prompting 0.715 0.664
SFT (Ours) 0.705 0.67